From the equation y+4x=16, write:
i) the equation in the slope intercept form.
ii) the gradient of the line.
iii) the y-intercept of the line.
Thank YOu soooo muchh:)

Answer by Scorinitron(19)

Y+4=16
(0)+4x=16
0+4x=16
4x=16
4x/4=16/4
x=4 (x intercept)
______________________________
y+4(0)=16
y+0=16
y=16 (y-intercept)
_____________________________
Slope intercept form is
Y=mx+b
y=-4x+16 (Slope Intercept Form)
_______________________________________
Gradient is another way of saying slope.
-4 (slope)

Your Answers
The Slope Intercept Form = y=-4x+16
Y-Intercept = 16
The Slope (gradient) of the line = -4